# pnasresearcharticle

### v1

@@ Abstract
Hydrogels are crosslinked polymer networks with high water content widely utilized in biomedical applications. Injectable depot-forming hydrogels for drug delivery offer sustained release of biomolecular cargo, leveraging controlled macromolecular diffusion within dynamic networks to promote extended immunity and favorable immunological responses. However, mitigating burst release while maintaining sustained-release properties remains a challenge due to network disruption during needle extrusion, leading to excessive short-term cargo release and local toxicity. While controlling both short- and long-term release profiles would mitigate this issue, it remains a persistent challenge. Here, we present a hydrogel formulation strategy for self-regulating, post-injection network evolution enabled by exchange of functional molecules with physiological environment, achieving controlled cargo release over both short and long durations. By tuning network viscoelasticity, we optimized rheological properties to minimize burst release while preserving sustained-release profiles. Using model protein cargo, in vitro and in vivo studies demonstrated reduced burst release and comparable sustained-release to unmodified control formulations. These findings highlight a novel strategy for precise cargo delivery in biomedical applications.
keywords:
Hydrogels |||| Cargo diffusion |||| Drug delivery |||| Rheology
[…]
This manuscript was compiled on March 19, 2025 www.pnas.org/cgi/doi/10.1073/pnas.XXXXXXXXXX
7]3
[…]
Hydrogels are crosslinked polymer networks with high water content, known for their biocompatibility and in vivo tolerability. Their highly tunable mechanochemical properties make them versatile materials for diverse applications (1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12). In the biomedical field, hydrogels have garnered significant attention, leading to discoveries of novel properties and applications (1, 3, 5, 6, 7, 9, 10, 11). Injectable, depot-forming hydrogels represent a critical subfield, enabling minimally invasive, sustained drug delivery. Using the slower diffusion of macromolecules, including proteins and other biologics, from dynamic hydrogel networks, these systems achieve prolonged therapeutic release, often far exceeding the durations of intravenous administration (cf. Fig. 1(a)), and elicit favorable immunological responses compared to bolus injections, underscoring their potential in advanced therapeutic applications (7, 13, 14, 15, 9, 6, 10, 3, 4, 16, 11).
A significant body of research focuses on developing hydrogels that prolong cargo retention in injected depots by tuning rheological properties (stiffness, relaxation time, yielding, and creep behavior) that control injection, depot persistence, and mass transport. The goal is to limit erosive cargo release and engineer a slow diffusive release rate (3, 13, 10, 17, 18, 19, 7). A typical sustained release profile from injectable hydrogel depots is shown in Fig. 1(a) (red). Physically crosslinked hydrogels with shear-thinning and self-healing properties enable easy injection and depot formation. Once injected, the release rates and cargo retention depend on the network structure, governed by its linear viscoelastic rheological properties (2, 3, 4, 20, 21, 5, 22, 23, 24, 16, 11). Increasing the stiffness of hydrogel networks (plateau shear modulus $G_{0}$) decreases the cargo release rate by reducing the effective mesh size ($\xi$), which impedes solute diffusion and lowers the diffusion coefficient compared to its Stokes-Einstein diffusivity ($\mathbb{D}/\mathbb{D}_{0}$). This relationship is supported by naïve network theory, which predicts $\xi\sim G_{0}^{-1/3}$, a trend confirmed by more advanced models (25, 26, 20, 19, 20, 13, 17, 23, 26, 16, 19).
The focus has primarily been on prolonging the duration of release by retaining cargo within the depot, rather than tailoring release profiles to control both short-term burst release and sustained release over longer periods (7, 6, 3, 10, 16, 1, 11). Despite excellent sustained release characteristics, most injectable hydrogel platforms exhibit burst release, leading to rapid absorption and therapeutically harmful local concentrations of drugs or antibodies, and high peak plasma concentrations, which may cause adverse effects (27, 28). Furthermore, burst release during the first few hours post-injection, as the hydrogel recovers its stiffness, reduces the duration over which the release remains within physiologically effective levels. The key challenge lies in minimizing burst release while maintaining appropriate sustained release profiles. Decoupling these two temporal regimes of mass transport to separately engineer each release phase presents significant rheological challenges.
@@
Our observations indicate that the dynamics of hydrogel networks, specifically the lifetimes and kinetics of the dynamic crosslinks that govern the self-diffusion of the network, play a key role in cargo diffusion, beyond the influence of topology (e.g., the average mesh size related to stiffness) (cf. Fig. 2). Although this phenomenon has been observed in previous studies (23, 24, 22), it remains underexplored. We propose that hydrogel structure recovery post-injection, during which crosslinks dissociate due to large strains in the narrow flow geometry, correlates with relaxation behavior: networks that relax stress rapidly also recover their structural integrity faster by reforming crosslinks. Hydrogels with slow crosslink kinetics exhibit favorable sustained release characteristics as a result of reduced cargo diffusion from their fully recovered structure. However, their slower recovery post-injection leads to excessive burst release, as loosely connected networks fail to arrest rapid diffusive or advective solute release driven by steep concentration gradients in cargo-free environments, such as intramuscular or subcutaneous spaces. In contrast, hydrogels with fast crosslink kinetics recover structure quickly, mitigating burst release by resisting mass transport. However, these materials often exhibit higher diffusion coefficients, compromising sustained release compared to slower relaxing networks.
@@
Results and Discussion
@@
Hydrogels with Modular Transport Properties: Linear Rheology and Diffusivity
The timescale of network relaxation, directly related to crosslinking kinetics, provides an additional degree of freedom in tuning cargo diffusivity within hydrogel meshes (22, 29). For networks with similar stiffness, those with distinct stress relaxation times exhibit different solute diffusivity, as observed in the hydrogels used in this study (cf. Fig. 2). This property is leveraged to design materials with rheological properties that evolve between fast and slow crosslink kinetics, controlling solute diffusion while maintaining similar stiffness. To achieve this, small molecules, such as specific ions or surfactants, are incorporated to modulate crosslinking and promote faster network relaxation. While this may enhance burst release, a slow diffusion of these molecules from the network allows the structure to gradually regain slower crosslinking kinetics, lowering the diffusion coefficient and prolonging cargo retention for sustained release.
To demonstrate that networks with similar stiffness but different relaxation times exhibit varying cargo diffusivity (Fig. 2(a, c)), we formulated a model transient hydrogel network (Fig. 1(c)). The system consisted of poly(vinyl alcohol) chains covalently crosslinked by borate ions. Borate- and boronic ester-crosslinked hydrogels are associative networks, where crosslinking kinetics directly influence relaxation behavior (30, 23, 31). The relaxation time and dissociation rate constant are related as $\tau_{\rm R}\simeq k_{\rm d}^{-1}$, as stress relaxation occurs via crosslink dissociation and reformation. Slowing dissociation kinetics significantly delays network relaxation. Modulating $p$H above the boric acid $pK_{\rm a}\simeq 9$ alters the crosslink dissociation kinetics, slowing relaxation without changing network stiffness, which remains governed by the average molecular weight between crosslinks. This is due to the $p$H-induced shift in association ($k_{\rm a}$) and dissociation ($k_{\rm d}$) rate constants, maintaining a constant equilibrium constant $K_{\rm eq}=k_{\rm a}/k_{\rm d}$, which is correlated with $G_{0}$. This orthogonal control over stiffness and relaxation time allows us to isolate the effect of network relaxation on cargo diffusion and formulate materials with comparable stiffness but different relaxation dynamics (cf. Fig. 2(a)).
The effect of network topology and dynamics on cargo release from hydrogels was bench-marked using in vitro release assays (13). We injected two PVA-Borax hydrogels with different relaxation times but similar stiffness ($G_{0}\simeq 675\leavevmode\nobreak\ {\rm Pa}$ and $735\leavevmode\nobreak\ {\rm Pa}$), into capillary tubes with buffers of matching pH and hydroxyl ion concentrations. This setup ensures a dynamic equilibrium between the hydrogel depot and the buffer interface, allowing only slow diffusive release of protein cargo. Cumulative release data ($M_{t}$) over 14 days, relative to the total cargo ($M_{0}$), is shown in Fig.3(c). To quantify these differences, we calculated the relative areas under the release profiles for both burst and sustained regimes (cf. Fig. 3(d)), such that $\alpha_{\rm burst}/\alpha_{\rm tot}$ and $\alpha_{\rm sust}/\alpha_{\rm tot}$, where $\alpha_{T_{1}T_{2}}=\sum_{n=T_{1}}^{T_{2}}(M_{t})_{n}t_{n}$ is the area between times $T_{1}$ and $T_{2}$, and $\alpha_{\rm tot}=\sum_{n=1}^{\rm end}(M_{t})_{n}t_{n}$ is the total area. For burst release, $T_{1}=0$ and $T_{2}=1\leavevmode\nobreak\ {\rm day}$; for sustained release, $T_{1}=2\leavevmode\nobreak\ {\rm days}$ and $T_{2}=14\leavevmode\nobreak\ {\rm days}$. As hypothesized, PVAB1/buffer1 releases less cargo in the burst phase due to its faster network relaxation, but more cargo over longer durations compared to PVAB3/buffer3. If PVAB1/buffer1 were functionalized to evolve into a slower relaxing network, we could potentially achieve both suppressed burst release and improved sustained release (vide infra).
In Vitro Cargo Release of Evolving Hydrogels
@@
In Vivo Pharmacokinetics of Evolving Hydrogels with Therapeutic Cargo in Mice
@@
As an ultimate test of applicability in a translational setting, we used the surfactant-modified Sangelose hydrogels to obtain in vivo pharmacokinetics (PK) data (Fig. 5). We assessed the serum protein concentration over time following injection of Sangelose hydrogels loaded with hIgG into mice, and the data were analyzed using a two-compartment pharmacokinetics model to describe the release kinetics (details in the SI). Both groups exhibited typical pharmacokinetics profiles with an initial increase in serum protein concentration followed by a decline. S3Tw0 reached a peak concentration ($c_{\rm max}$) of $282.27\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ mL}^{-1}$ at $48\leavevmode\nobreak\ {\rm h}$ ($t_{\rm max}$), while S3Tw2 reached a peak of $260.69\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ mL}^{-1}$ at $96\leavevmode\nobreak\ {\rm h}$. It is clear that S3Tw0 demonstrated a faster release profile, with a more pronounced burst release, as indicated by the earlier $t_{\rm max}$ and higher $c_{\rm max}$. In contrast, S3Tw2 exhibited a more sustained release, with slower decline in serum concentrations, as evidenced by its larger area under the curve (AUC) of $55,181.69\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ h\leavevmode\nobreak\ mL}^{-1}$, compared to $50,752.97\leavevmode\nobreak\ \mu{\rm g\leavevmode\nobreak\ h\leavevmode\nobreak\ mL}^{-1}$ for S3Tw0. Fits of the PK data (cf. Fig. 5(c)) to a two-compartment model also yield the half-lives for (i) depot/cargo release into the subcutaneous space and subsequent absorption into the bloodstream, $\tau_{\rm r}$ (compartment 1) and (ii) elimination from the body, $\tau_{\rm e}$ (compartment 2) (details in the SI). From the model fits, $\tau_{\rm r}$ was obtained to be $3.19\pm 1.27$ and $6.02\pm 1.04\leavevmode\nobreak\ {\rm d}$ for S3Tw0 and S3Tw2 respectively (cf. Fig. 5(e), whereas $\tau_{\rm e}$ was found to be $1.30\pm 0.40$ and $1.01\pm 0.16\leavevmode\nobreak\ {\rm d}$ respectively. Elimination kinetics are typically assumed to be independent of administration method and vehicle, and thus the values being essentially the same across both groups with very similar hydrogel structure, chemistry, and biological compatibility are consistent with previous literature (13, 35). The only significant difference between the two groups was observed in the release half-lives, where S3Tw2 released cargo significantly slower than S3Tw0. These findings highlight the improved burst-release performance of Sangelose functionalized with tween 80, achieving slower overall release while maintaining sustained cargo delivery.
Discussion
@@
This study highlights the potential of injectable dynamic hydrogels to intrinsically evolve their transport properties post-injection within subcutaneous depots, offering a novel approach to mitigating burst release while modulating cargo release rates. By engineering the magnitude and direction of changes in rheological and mass transport properties through functional molecules that influence crosslinking kinetics, we achieve hybrid release profiles seamlessly transitioning between burst and sustained release. This self-regulated evolution addresses a critical challenge in injectable hydrogels, mitigating sudden, excessive cargo release while maintaining effective sustained delivery. Such doubly-controlled release profiles are beneficial for a diverse array of therapeutics deliveries: minimal burst release is strongly desirable for delivery of therapeutics with strong toxicities, e.g. GLP-1 agonists in diabetes treatment (36); slow, prolonged release is essential for ultra-long administration of precious drug cargo for treating various infectious diseases (13). Decoupling and independently controlling these release regimes overcomes the inherent constraints of stable crosslinked networks. Modular control over network relaxation, recovery, and solute diffusivity offers a blueprint for designing hydrogels tailored to the complex demands of drug delivery and other biomedical applications. Our findings establish a foundation for next-generation soft materials capable of dynamically adapting their structure and function to physiological conditions. This advancement enables more precise therapeutic delivery and bridges the gap between material design and clinical application. Future work can expand on these insights to explore broader applications and refine control mechanisms for complex release profiles, unlocking the full potential of injectable hydrogels in diverse biomedical settings.
Conflicts of Interest
[…]

### v2

# Evolving transport properties of dynamic hydrogels enable self-tuning of short- and long-term cargo delivery

@@ Abstract
Keywords: Hydrogels, Cargo diffusion, Drug delivery, Rheology
1. Introduction
Hydrogels are crosslinked polymer networks with high water content, known for their biocompatibility and in vivo tolerability. Their highly tunable mechanochemical properties make them versatile materials for diverse applications [9, 1, 20, 2, 37, 18, 33, 30, 7, 26, 19, 11]. In the biomedical field, hydrogels have garnered significant attention, leading to discoveries of novel properties and applications [9, 20, 37, 18, 33, 7, 26, 19]. Injectable, depot-forming hydrogels represent a critical subfield, enabling minimally invasive, sustained drug delivery. Using the slower diffusion of macromolecules, including proteins and other biologics, from dynamic hydrogel networks, these systems achieve prolonged therapeutic release, often far exceeding the durations of intravenous administration (cf. Fig. 1(a)), and elicit favorable immunological responses compared to bolus injections, underscoring their potential in advanced therapeutic applications [33, 17, 12, 13, 7, 18, 26, 20, 2, 29, 19].
@@ 1. Introduction
A significant body of research focuses on developing hydrogels that prolong cargo retention in injected depots by tuning rheological properties (stiffness, relaxation time, yielding, and creep behavior) that control injection, depot persistence, and mass transport. The goal is to limit erosive cargo release and engineer a slow diffusive release rate [20, 17, 26, 32, 16, 28, 33]. A typical sustained release profile from injectable hydrogel depots is shown in Fig. 1(a) (red). Physically crosslinked hydrogels with shear-thinning and self-healing properties enable easy injection and depot formation. Once injected, the release rates and cargo retention depend on the network structure, governed by its linear viscoelastic rheological properties [1, 20, 2, 5, 15, 37, 6, 22, 23, 29, 19]. Increasing the stiffness of hydrogel networks (plateau shear modulus $G_{0}$) decreases the cargo release rate by reducing the effective mesh size ($\xi$), which impedes solute diffusion and lowers the diffusion coefficient compared to its Stokes-Einstein diffusivity ($\mathbb{D}/\mathbb{D}_{0}$). This relationship is supported by naïve network theory, which predicts $\xi\sim G_{0}^{-1/3}$, a trend confirmed by more advanced models [31, 27, 5, 28, 17, 32, 22, 29].
The focus has primarily been on prolonging the duration of release by retaining cargo within the depot, rather than tailoring release profiles to control both short-term burst release and sustained release over longer periods [33, 18, 20, 26, 29, 9, 19]. Despite excellent sustained release characteristics, most injectable hydrogel platforms exhibit burst release, leading to rapid absorption and therapeutically harmful local concentrations of drugs or antibodies, and high peak plasma concentrations, which may cause adverse effects [8, 14]. Furthermore, burst release during the first few hours post-injection, as the hydrogel recovers its stiffness, reduces the duration over which the release remains within physiologically effective levels. The key challenge lies in minimizing burst release while maintaining appropriate sustained release profiles. Decoupling these two temporal regimes of mass transport to separately engineer each release phase presents significant rheological challenges.
Our observations indicate that the dynamics of hydrogel networks, specifically the lifetimes and kinetics of the dynamic crosslinks that govern the self-diffusion of the network, play a key role in cargo diffusion, beyond the influence of topology (e.g., the average mesh size related to stiffness) (cf. Fig. 2). Although this phenomenon has been observed in previous studies [22, 23, 6], it remains underexplored. We propose that hydrogel structure recovery post-injection, during which crosslinks dissociate due to large strains in the narrow flow geometry, correlates with relaxation behavior: networks that relax stress rapidly also recover their structural integrity faster by reforming crosslinks. Hydrogels with slow crosslink kinetics exhibit favorable sustained release characteristics as a result of reduced cargo diffusion from their fully recovered structure. However, their slower recovery post-injection leads to excessive burst release, as loosely connected networks fail to arrest rapid diffusive or advective solute release driven by steep concentration gradients in cargo-free environments, such as intramuscular or subcutaneous spaces. In contrast, hydrogels with fast crosslink kinetics recover structure quickly, mitigating burst release by resisting mass transport. However, these materials often exhibit higher diffusion coefficients, compromising sustained release compared to slower relaxing networks.
@@
2. Results
2.1. Hydrogels with Modular Transport Properties: Rheology and Diffusivity
The timescale of network relaxation, directly related to crosslinking kinetics, provides an additional degree of freedom in tuning cargo diffusivity within hydrogel meshes [6, 3]. For networks with similar stiffness, those with distinct stress relaxation times exhibit different solute diffusivity, as observed in the hydrogels used in this study (cf. Fig. 2). This property is leveraged to design materials with rheological properties that evolve between fast and slow crosslink kinetics, controlling solute diffusion while maintaining similar stiffness. To achieve this, small molecules, such as specific ions or surfactants, are incorporated to modulate crosslinking and promote faster network relaxation. While this may enhance burst release, a slow diffusion of these molecules from the network allows the structure to gradually regain slower crosslinking kinetics, lowering the diffusion coefficient and prolonging cargo retention for sustained release.
@@ 2.1. Hydrogels with Modular Transport Properties: Rheology and Diffusivity
To demonstrate that networks with similar stiffness but different relaxation times exhibit varying cargo diffusivity (Fig. 2(a, c)), we formulated a model transient hydrogel network (Fig. 1(c)). The system consisted of poly(vinyl alcohol) chains covalently crosslinked by borate ions. Borate- and boronic ester-crosslinked hydrogels are associative networks, where crosslinking kinetics directly influence relaxation behavior [24, 22, 36]. The relaxation time and dissociation rate constant are related as $\tau_{\rm R}\simeq k_{\rm d}^{-1}$, as stress relaxation occurs via crosslink dissociation and reformation. Slowing dissociation kinetics significantly delays network relaxation. Modulating $p$H above the boric acid $pK_{\rm a}\simeq 9$ alters the crosslink dissociation kinetics, slowing relaxation without changing network stiffness, which remains governed by the average molecular weight between crosslinks. This is due to the $p$H-induced shift in association ($k_{\rm a}$) and dissociation ($k_{\rm d}$) rate constants, maintaining a constant equilibrium constant $K_{\rm eq}=k_{\rm a}/k_{\rm d}$, which is correlated with $G_{0}$. This orthogonal control over stiffness and relaxation time allows us to isolate the effect of network relaxation on cargo diffusion and formulate materials with comparable stiffness but different relaxation dynamics (cf. Fig. 2(a)).
The effect of network topology and dynamics on cargo release from hydrogels was bench-marked using in vitro release assays [17]. We injected two PVA-Borax hydrogels with different relaxation times but similar stiffness ($G_{0}\simeq 675~{}{\rm Pa}$ and $735~{}{\rm Pa}$), into capillary tubes with buffers of matching pH and hydroxyl ion concentrations. This setup ensures a dynamic equilibrium between the hydrogel depot and the buffer interface, allowing only slow diffusive release of protein cargo. Cumulative release data ($M_{t}$) over 14 days, relative to the total cargo ($M_{0}$), is shown in Fig.3(c). To quantify these differences, we calculated the relative areas under the release profiles for both burst and sustained regimes (cf. Fig. 3(d)), such that $\alpha_{\rm burst}/\alpha_{\rm tot}$ and $\alpha_{\rm sust}/\alpha_{\rm tot}$, where $\alpha_{T_{1}T_{2}}=\sum_{n=T_{1}}^{T_{2}}(M_{t})_{n}t_{n}$ is the area between times $T_{1}$ and $T_{2}$, and $\alpha_{\rm tot}=\sum_{n=1}^{\rm end}(M_{t})_{n}t_{n}$ is the total area. For burst release, $T_{1}=0$ and $T_{2}=1~{}{\rm day}$; for sustained release, $T_{1}=2~{}{\rm days}$ and $T_{2}=14~{}{\rm days}$. As hypothesized, PVAB1/buffer1 releases less cargo in the burst phase due to its faster network relaxation, but more cargo over longer durations compared to PVAB3/buffer3. If PVAB1/buffer1 were functionalized to evolve into a slower relaxing network, we could potentially achieve both suppressed burst release and improved sustained release (vide infra).
@@
2.2. In Vitro Cargo Release of Evolving Hydrogels
[…]
@@ 2.2. In Vitro Cargo Release of Evolving Hydrogels
2.3. In Vivo Pharmacokinetics of Evolving Hydrogels with Therapeutic Cargo in Mice
@@
As an ultimate test of applicability in a translational setting, we used the surfactant-modified Sangelose hydrogels to obtain in vivo pharmacokinetics (PK) data (Fig. 5). We assessed the serum protein concentration over time following injection of Sangelose hydrogels loaded with hIgG into mice, and the data were analyzed using a two-compartment pharmacokinetics model to describe the release kinetics (details in the SI). Both groups exhibited typical pharmacokinetics profiles with an initial increase in serum protein concentration followed by a decline. S3Tw0 reached a peak concentration ($c_{\rm max}$) of $282.27~{}\mu{\rm g~{}mL}^{-1}$ at $48~{}{\rm h}$ ($t_{\rm max}$), while S3Tw2 reached a peak of $260.69~{}\mu{\rm g~{}mL}^{-1}$ at $96~{}{\rm h}$. It is clear that S3Tw0 demonstrated a faster release profile, with a more pronounced burst release, as indicated by the earlier $t_{\rm max}$ and higher $c_{\rm max}$. In contrast, S3Tw2 exhibited a more sustained release, with slower decline in serum concentrations, as evidenced by its larger area under the curve (AUC) of $55,181.69~{}\mu{\rm g~{}h~{}mL}^{-1}$, compared to $50,752.97~{}\mu{\rm g~{}h~{}mL}^{-1}$ for S3Tw0. Fits of the PK data (cf. Fig. 5(c)) to a two-compartment model also yield the half-lives for (i) depot/cargo release into the subcutaneous space and subsequent absorption into the bloodstream, $\tau_{\rm r}$ (compartment 1) and (ii) elimination from the body, $\tau_{\rm e}$ (compartment 2) (details in the SI). From the model fits, $\tau_{\rm r}$ was obtained to be $3.19\pm 1.27$ and $6.02\pm 1.04~{}{\rm d}$ for S3Tw0 and S3Tw2 respectively (cf. Fig. 5(e), whereas $\tau_{\rm e}$ was found to be $1.30\pm 0.40$ and $1.01\pm 0.16~{}{\rm d}$ respectively. Elimination kinetics are typically assumed to be independent of administration method and vehicle, and thus the values being essentially the same across both groups with very similar hydrogel structure, chemistry, and biological compatibility are consistent with previous literature [17, 21]. The only significant difference between the two groups was observed in the release half-lives, where S3Tw2 released cargo significantly slower than S3Tw0. These findings highlight the improved burst-release performance of Sangelose functionalized with tween 80, achieving slower overall release while maintaining sustained cargo delivery.
3. Discussion
@@
This study highlights the potential of injectable dynamic hydrogels to intrinsically evolve their transport properties post-injection within subcutaneous depots, offering a novel approach to mitigating burst release while modulating cargo release rates. By engineering the magnitude and direction of changes in rheological and mass transport properties through functional molecules that influence crosslinking kinetics, we achieve hybrid release profiles seamlessly transitioning between burst and sustained release. This self-regulated evolution addresses a critical challenge in injectable hydrogels, mitigating sudden, excessive cargo release while maintaining effective sustained delivery. Such doubly-controlled release profiles are beneficial for a diverse array of therapeutics deliveries: minimal burst release is strongly desirable for delivery of therapeutics with strong toxicities, e.g. GLP-1 agonists in diabetes treatment [10]; slow, prolonged release is essential for ultra-long administration of precious drug cargo for treating various infectious diseases [17]. Decoupling and independently controlling these release regimes overcomes the inherent constraints of stable crosslinked networks. Modular control over network relaxation, recovery, and solute diffusivity offers a blueprint for designing hydrogels tailored to the complex demands of drug delivery and other biomedical applications. Our findings establish a foundation for next-generation soft materials capable of dynamically adapting their structure and function to physiological conditions. This advancement enables more precise therapeutic delivery and bridges the gap between material design and clinical application. Future work can expand on these insights to explore broader applications and refine control mechanisms for complex release profiles, unlocking the full potential of injectable hydrogels in diverse biomedical settings.
Methods